# WHAM: Reconstructing World-grounded Humans with Accurate 3D Motion

### v1

@@ A.1 Synthetic data generation
The goal of generating a contact label is to train the motion decoder to detect the foot-ground contact accurately. Previous work [60] uses both the velocity and height of the feet to generate contact labels. However, in order to generalize our approach to arbitrary ground conditions such as slopes or stairs, we only use foot velocity to compute the ground truth contact labels, similar to TransPose [55]. We use heel and toe vertices of each foot to define foot-ground contact and compute the probability as below:
@@
We set the threshold velocity $v_{t}=1cm/\text{frame}$ and the coefficient $\alpha=5$.
@@
$\gamma_{r}^{(0)}$ $\sim$ $\mathcal{N}(0^{\circ},5^{\circ}),$
$\gamma_{p}^{(0)}$ $\sim$ $\mathcal{N}(5^{\circ},22.5^{\circ}).$
@@
Subsequently, we sample the initial camera translation, using a mix of unfiorm and normal distributions, to capture the ground-truth 3D pose in the camera coordinates
$T_{z}^{(0)}$ $\sim$ $\mathcal{U}(2m,12m)-\big{(}R^{(0)}t^{(0)}\big{)}_{z},$
$T_{x}^{(0)}$ $\sim$ $\mathcal{N}(0,0.25d)~{}d-\big{(}R^{(0)}t^{(0)}\big{)}_{x},$
$T_{y}^{(0)}$ $\sim$ $\mathcal{N}(0,0.25d)~{}d-\big{(}R^{(0)}t^{(0)}\big{)}_{y}.$
@@
$\Delta\gamma_{y}$ $\sim$ $\mathcal{N}(0^{\circ},45^{\circ}),$
$\Delta\gamma_{r},\Delta\gamma_{p}$ $\sim$ $\mathcal{N}(0^{\circ},22.5^{\circ}),$
$\Delta T_{x},\Delta T_{y},\Delta T_{z}$ $\sim$ $\mathcal{N}(0m,1m).$
@@ A.2 Datasets
3DPW
Methods   PA-MPJPE MPJPE PVE Accel
w/o Neural State Initialization   37.8 64.5 76.3 6.9
WHAM   37.8 60.8 72.5 6.8
Table 4: Quantitative analysis of neural state initialization on the 3DPW [49] dataset.
A.3 Neural-network Initialization
@@
Uni-directional RNNs introduce the challenge of differing learning objectives between the initial frames and subsequent ones due to the initialization state. Specifically, in traditional RNNs, the initialization state is typically padded with zeros, resulting in the first frame primarily relying on the input signal. In contrast, the subsequent frames are trained to capitalize on both the input signal and information transferred from the past. To resolve the disparity in learning objectives, we use a neural initialization network, as proposed by [56], to predict $h_{0,E}$ and $h_{0,D}$ from the 0-th frame pose, instead of using zero-padding. During the training, we use pseudo-ground-truth 3D pose [35] for the video datasets that do not have the SMPL parameter annotation [11, 33, 15]. Note that we do not supervise our network on the pseudo labels. At test time, we use the pose and shape predicted by a single-frame regressor as the initial state. As shown in Table 4, we observe that the use of neural state initialization increases the performance in 3D human pose and shape estimation.
@@
A.4 Run-time cost

### v2

@@ Appendix A Synthetic data generation
The goal of generating a contact label is to train the motion decoder to detect the foot-ground contact accurately. Previous work [61] uses both the velocity and height of the feet to generate contact labels. However, in order to generalize our approach to arbitrary ground conditions such as slopes or stairs, we only use foot velocity to compute the ground truth contact labels, similar to TransPose [56]. We use the heel and toe vertices of each foot to define foot-ground contact and compute the probability as below:
@@
We set the threshold velocity $v_{t}=1\text{cm}/\text{frame}$ and the coefficient $\alpha=5$.
@@
$\gamma_{r}^{(0)}$ $\sim$ $\mathcal{N}(0^{\circ},(5^{\circ})^{2}),$
$\gamma_{p}^{(0)}$ $\sim$ $\mathcal{N}(5^{\circ},(22.5^{\circ})^{2}).$
@@
Subsequently, we sample the initial camera translation, using a mix of uniform and normal distributions, to capture the ground-truth 3D pose in the camera coordinates:
$T_{z}^{(0)}$ $\sim$ $\mathcal{U}(2m,(12\rm{m})^{2})-\big{(}R^{(0)}t^{(0)}\big{)}_{z},$
$T_{x}^{(0)}$ $\sim$ $\mathcal{N}(0,0.25^{2})~{}d-\big{(}R^{(0)}t^{(0)}\big{)}_{x},$
$T_{y}^{(0)}$ $\sim$ $\mathcal{N}(0,0.25^{2})~{}d-\big{(}R^{(0)}t^{(0)}\big{)}_{y}.$
@@
$\Delta\gamma_{y}$ $\sim$ $\mathcal{N}(0^{\circ},(45^{\circ})^{2}),$
$\Delta\gamma_{r},\Delta\gamma_{p}$ $\sim$ $\mathcal{N}(0^{\circ},(22.5^{\circ})^{2}),$
$\Delta T_{x},\Delta T_{y},\Delta T_{z}$ $\sim$ $\mathcal{N}(0m,(1m)^{2}).$
@@ Appendix B Datasets
Appendix C Losses
The loss functions of WHAM can be categorized into two groups; motion reconstruction and trajectory reconstruction:
$L_{\rm{total}}$ $=$ $\overbrace{L_{\rm{smpl}}+L_{\rm{verts}}+L_{\rm{3D}}+L_{\rm{2D}}+L_{\rm{casc}}}^{\text{motion reconstruction}}+$
$\overbrace{L_{\rm{root}}+L_{\rm{contact}}+L_{\omega}+L_{\rm{cam}}+L_{\rm{fs}}}^{\text{trajectory reconstruction}}.$
To be specific, we supervise WHAM directly on SMPL parameters and vertices when we have the labels, $\theta^{*}$, $\beta^{*}$ and $V^{*}$:
$L_{\rm{smpl}}$ $=$ $\sum_{t=0}^{T}\left(\lambda_{\theta}||\theta^{(t)}-\theta^{*(t)}||_{2}+\lambda_{\beta}||\beta^{(t)}-\beta^{*(t)}||_{2}\right),$
$L_{\rm{verts}}$ $=$ $\sum_{t=0}^{T}\lambda_{\rm{verts}}||V^{(t)}-V^{*(t)}||_{1}.$
When the 3D keypoints annotations $(x_{3D}^{*})$ are available, we follow the standard MSE loss on 3D keypoints directly regressed from the Motion Encoder $(x_{3D})$ and extracted from the predicted SMPL mesh $(\hat{x}_{\rm{3D}})$:
$L_{\rm{3D}}=\sum_{t=0}^{T}\lambda_{\rm{3D}}\left(||x_{\rm{3D}}^{(t)}-x_{\rm{3D}}^{*(t)}||_{2}+||\hat{x}_{\rm{3D}}^{(t)}-x_{\rm{3D}}^{*(t)}||_{2}\right).$
Since WHAM reconstructs full-body mesh in a cascaded manner, we use cascade loss $L_{\textrm{casc}}$ to supervise the difference between $x_{\rm{3D}}$ and $\hat{x}_{\rm{3D}}$:
$L_{\rm{casc}}=\sum_{t=0}^{T}\lambda_{\rm{casc}}||x_{\rm{3D}}^{(t)}-\hat{x}_{\rm{3D}}||_{2}.$
For 2D loss, we follow CLIFF [25] to compute full-perspective 2D reprojection loss when the label is available, while weak-perspective 2D loss is applied for InstaVariety dataset as its full-frame image is not available:
$L_{\rm{2D}}$ $=$ $\sum_{t=0}^{T}\lambda_{\rm{2D}}||\pi\left(\hat{x}_{\rm{3D}}^{(t)}\right)-x_{\rm{2D}}^{*(t)}||_{2},$
where $\pi(\cdot)$ denotes the operation of perspective projection.
We supervise root trajectory and foot contact probability directly using MSE loss when the ground truth data is available:
$L_{\rm{root}}$ $=$ $\sum_{t=0}^{T}\lambda_{\Gamma}\left(||\Gamma_{0}^{(t)}-\Gamma^{*(t)}||_{2}+||\Gamma^{(t)}-\Gamma^{*(t)}||_{2}\right)+$
$\sum_{t=0}^{T}\lambda_{v}\left(||v_{0}^{(t)}-v^{*(t)}||_{2}+||v^{(t)}-v^{*(t)}||_{2}\right),$
$L_{\rm{contact}}$ $=$ $\sum_{t=0}^{T}\lambda_{p}||p^{(t)}-p^{*(t)}||_{2}.$
The Camera’s pose can be expressed by the predicted root orientation in world and camera coordinates:
$R^{(t)}$ $=$ $\Gamma_{c}^{(t)}\left(\Gamma^{(t)}\right)^{\top},$
where $\Gamma_{c}$ is the predicted root orientation in the camera coordinate system. As a result, we can construct novel losses based on camera pose and angular velocity:
$L_{\omega}$ $=$ $\sum_{t=0}^{T}\lambda_{\omega}||\omega^{(t)}-\omega^{*(t)}||_{2},$
$L_{\rm{cam}}$ $=$ $\sum_{t=0}^{T}\lambda_{\rm{cam}}||R^{(t)}-R^{*(t)}||_{2},$
where $\omega$ is the reconstructed camera angular velocities from $R$.
Finally, we add foot sliding loss using the foot velocity during the contact:
$L_{\rm{fs}}$ $=$ $\sum_{t=0}^{T}\lambda_{\rm{fs}}||p^{*(t)}\odot v_{f}^{(t)}||_{2}.$
Here, $v_{f}^{(t)}$ is the world-coordinate foot velocity, and $\odot$ denotes the masking operation of foot contact based on the ground truth contact probability $p^{*(t)}$.
Appendix D Neural-network initialization
@@
Uni-directional RNNs introduce the challenge of differing learning objectives between the initial frames and subsequent ones due to the initialization state. Specifically, in traditional RNNs, the initialization state is typically padded with zeros, resulting in the first frame primarily relying on the input signal. In contrast, the subsequent frames are trained to capitalize on both the input signal and information transferred from the past. To resolve the disparity in learning objectives, we use a neural initialization network, as proposed by [57], to predict $h_{0,E}$ and $h_{0,D}$ from the 0-th frame pose, instead of using zero-padding. During the training, we use pseudo-ground-truth 3D pose [35] for the video datasets that do not have the SMPL parameter annotation [11, 33, 15]. Note that we do not supervise our network on the pseudo labels. At test time, we use the pose and shape predicted by a single-frame regressor as the initial state.
Figure 7: Qualitative comparison between WHAM and after removal of contact-aware trajectory refinement (w/o traj. ref.).
Appendix E Qualitative evaluation of contact estimation
In the main paper, we show the quantitative effect of the contact-aware trajectory refinement unit. To supplement this with more intuitive visualization, we provide qualitative results in Fig. 7. As shown in the figure, WHAM without the trajectory refinement unit induces significant foot sliding during the contact. Conversely, WHAM reconstructs a more feasible trajectory based on its contact estimation.
Appendix F Run-time cost
@@
Figure 8: Failure cases of WHAM in global motion estimation.
Appendix G Discussions of limitation
In Fig. 8, we demonstrate exemplar cases when WHAM fails to capture the global motion of the person. Since the datasets we used for training WHAM on global trajectory estimation do not contain people riding skateboards or bicycles, WHAM does not capture the global trajectory in these scenarios. Furthermore, our contact estimation only applies to the feet, thus, WHAM fails to capture ground contact of the human body other than the feet, resulting in physically infeasible body support (floating hands) or sliding of the contact points. More information on human-object interactions can be used to resolve this issue. While we employ random masking as part of our data synthesis process, our generating approach mainly assumes the scenario where the full body is within the field of view. This can be addressed with additional augmentation during training (cf. [18, 26]).